which equation of the slopes of two nonvertical perpendicular lines is
A.y=-2/5x+3
B.y=-5/2x+2
C.y=-5/2x+3
D.y=3x+3
E.y=5/2x+6

Answer by stanbon(75887)

A.y=-2/5x+3
B.y=-5/2x+2
C.y=-5/2x+3
D.y=3x+3
E.y=5/2x+6
------------
A and E because the product of their slopes is -1.
(-2/5)(5/2)=-1
Cheers
Stan H.